a sample nitrogen gas undergoes an adiabatic expansion. The initial pressure and volume of the gas is 4.00 atm and 14.0 L. The final volume is 30.0L. If there are 2 moles of gas, determine the following


1
Expert's answer
2021-09-27T18:23:09-0400

Solution.

"P_i=4.00 atm;"

"V_i=14.0L;"

"V_f=30.0L;"

"\gamma=1.4;"

"\dfrac{P_f}{P_i}=(\dfrac{V_i}{V_f})^\gamma\implies P_f=P_i(\dfrac{V_i}{V_f})^\gamma;"

"P_f=4.00\sdot(\dfrac{14.0}{30.0})^{1.4}=1.38 atm;"

Answer: "P_f=1.38 atm."